# An Approximation Algorithm for Monotone Submodular Cost Allocation

### v1

@@ 2 Approximation algorithm
Since $M$ is divisible by k(k1)k(k-1), we can denote $M=k(k-1)\cdot m$ for some integer $m$. For $k$ positive integers $a_{1},\ldots,a_{k}\in[M]$, we say that an ordered tuple $(a_{1},\ldots,a_{k})$ covers $N$ if $\bigcup_{i=1}^{k}U_{i}^{a_{i}}=N$. We show that if there exist distinct $M\cdot 2/k=2m(k-1)$ ordered tuples covering $N$, then one of them corresponds to an integer solution to Mono-LP-Rel that achieves $k/2$-approximation.
@@
    each ordered tuple covers $N$.
Then, the integrality gap of Mono-LP-Rel is at most $k/2$.
@@
$\sum_{S\subseteq N,i\in[k]}y^{*}_{i}(S)f_{i}(S)=\sum_{i\in[k],j\in[M]}\dfrac{1}{M}f_{i}(U_{i}^{j})\geq\sum_{i\in[k],j\in[2m(k-1)]}\dfrac{1}{M}f_{i}(U_{i}^{a_{i}^{j}})$
$\geq\dfrac{2m(k-1)}{M}\cdot\min\left\{\sum_{i\in[k]}f_{i}(U_{i}^{a_{i}^{j}})\ \middle|\ j\in[2m(k-1)]\right\}$
$=\dfrac{2}{k}\cdot\min\left\{\sum_{i\in[k]}f_{i}(U_{i}^{a_{i}^{j}})\ \middle|\ j\in[2m(k-1)]\right\}.$ (2.2)
@@
$\sum_{S\subseteq N,i\in[k]}y^{*}_{i}(S)f_{i}(S)\geq\dfrac{2}{k}\sum_{i\in[k]}f_{i}(U_{i}^{a_{i}^{j^{*}}}).$
Hence, we have
@@
Since $\bigcup_{i=1}^{k}U_{i}^{a_{i}^{j^{*}}}=N$, there exists a partition $U_{1},\ldots,U_{k}$ of $N$ such that $U_{i}\subseteq U_{i}^{a_{i}^{j^{*}}}$ for every $i\in[k]$. Then, by (2.3) and the monotonicity of $f_{i}$, we have
$\dfrac{k}{2}\cdot\sum_{S\subseteq N,i\in[k]}y^{*}_{i}(S)f_{i}(S)\geq\sum_{i\in[k]}f_{i}(U_{i}^{a_{i}^{j^{*}}})\geq\sum_{i\in[k]}f_{i}(U_{i}).$
This implies that the integrality gap of Mono-LP-Rel is at most $k/2$. ∎
We are now ready to prove Theorem 1.1.
Theorem 1.1.
By Lemma 2.4, it suffices to show that there exist 2m(k1)2m(k-1) ordered tuples satisfying the conditions in Lemma 2.4. For each positive integer $a\in\mathbb{Z}_{>0}$, define $\mathrm{mod}(a)$ as follows:
[…]
For each $i\in[k]$ and $j\in[2m(k-1)]$, define $a_{i}^{j}$ as follows:
@@
We show that $(a_{i}^{j})_{i\in[k],j\in[2m(k-1)]}$ satisfies conditions (1)–(3) in Lemma 2.4. For condition (1), we show that $1\leq a_{i}^{j}\leq M$ for each $i\in[k]$ and $j\in[2m(k-1)]$. If $1\leq i\leq k-1$, then we have
@@
The proof of Theorem 1.1 provides the construction of a $k/2$-approximate integer solution to Mono-LP-Rel. We prove Theorem 1.2 by showing that such an integer solution can be constructed in polynomial time.
Theorem 1.2.
We can obtain a rational optimal solution $(y_{1}^{*},\ldots,y_{k}^{*})$ to Mono-LP-Rel whose supports $\mathcal{L}_{1},\ldots,\mathcal{L}_{k}$ are chains in polynomial time via the ellipsoid method [5] (see Appendix A for details). Then, we define $U_{i}^{j}$ for each $i\in[k]$ and $j\in[M]$, and $a_{i}^{j}$ for each $i\in[k]$ and $j\in[2m(k-1)]$ as in the proof of Theorem 1.1. If we can compute
$j^{*}:=\underset{j\in[2m(k-1)]}{\operatorname{argmin}}\sum_{i\in[k]}f_{i}(U_{i}^{a_{i}^{j}}),$
then any partition $U_{1},\ldots,U_{k}$ of $N$ such that $U_{i}\subseteq U_{i}^{a_{i}^{j^{*}}}$ for every $i\in[k]$ satisfies
$\sum_{i\in[k]}f_{i}(U_{i})\leq\dfrac{k}{2}\cdot\sum_{S\subseteq N,i\in[k]}y^{*}_{i}(S)f_{i}(S)\leq\dfrac{k}{2}\cdot\min\left\{\sum_{i=1}^{k}f_{i}(X_{i})\ \middle|\ X_{1},\ldots,X_{k}\ \mathrm{is\ a\ partition\ of\ }N\right\},$
as desired. Hence, it suffices to show that $j^{*}$ can be computed in polynomial time. To find $j^{*}$, we set $j_{1}:=1$. Then, for q=2,3,,q=2,3,..., we define $j_{q}$ as the minimum integer such that $j_{q}>j_{q-1}$ and there exists $i\in[k]$ with $U_{i}^{a_{i}^{j_{q}}}\neq U_{i}^{a_{i}^{j_{q-1}}}$. Recall that $n:=|N|$. Since for each $i\in[k]$, we have $U_{i}^{1}\supseteq U_{i}^{2}\supseteq\cdots\supseteq U_{i}^{M}$ and there are at most $|\mathcal{L}_{i}|\leq n$ distinct sets among $\{U_{i}^{j}\}_{j=1}^{M}$, we only need to define $j_{q}$ for $q=1,2,\ldots,Q$ with $Q\leq kn$. Since we have
$j^{*}=\underset{j\in\{j_{1},j_{2},\ldots,j_{Q}\}}{\operatorname{argmin}}\sum_{i\in[k]}f_{i}(U_{i}^{a_{i}^{j}}),$
we can find $j^{*}$ by computing the values of $\sum_{i\in[k]}f_{i}(U_{i}^{a_{i}^{j_{q}}})$ for all $q\in[Q]$. By the definition of $U_{i}^{j}$ and $a_{i}^{j}$, we can compute $j_{q}$ in polynomial time for each fixed $q$. Hence, since $Q\leq kn$, we can compute $j_{q}$ and $\sum_{i\in[k]}f_{i}(U_{i}^{a_{i}^{j_{q}}})$ for all $q\in[Q]$ in polynomial time. ∎
@@ 3 Lower bound on the integrality gap
In this section, we show that the integrality gap of Mono-LP-Rel is at least $\frac{k}{2}-\frac{k(k-1)}{\log n}$ when $\log n\geq k$. To this end, we explicitly construct monotone submodular functions $f_{1},\ldots,f_{k}$ such that the integrality gap of Mono-LP-Rel is at least $\frac{k}{2}-\frac{k(k-1)}{\log n}$.
@@
We now show that for the monotone submodular functions $f_{1},\ldots,f_{k}$ defined above, any optimal solution $(y_{1}^{*},\ldots,y_{k}^{*})$ to Mono-LP-Rel satisfies
@@
$\leq\min\left\{\sum_{i=1}^{k}f_{i}(X_{i})\ \middle|\ X_{1},\ldots,X_{k}\ \mathrm{is\ a\ partition\ of\ }N\right\}.$ (3.1)
The inequality (3) implies that the integrality gap of Mono-LP-Rel is at least $\frac{k}{2}-\frac{k(k-1)}{\log n}$, which completes the proof of Theorem 1.3.
To prove (3), we first give an upper bound on the value $\sum_{S\subseteq N,i\in[k]}y^{*}_{i}(S)f_{i}(S)$. To this end, we construct a feasible solution to Mono-LP-Rel. For each $i\in[k]$ and $j\in[pk-k+1]$, let $N_{i}^{j}:=\{e_{v}\in N\mid v\in\mathcal{V},v_{i}\geq j\}$. Note that $N_{i}^{1}=N_{i}$. In what follows, we will use $N_{i}^{1}$ instead of $N_{i}$ for convenience. Let $N^{c}:=N\setminus\{e_{v}\in N\mid v\in\mathcal{V}\}$. For each $i\in[k]$, define $z_{i}\in\mathbb{R}_{+}^{2^{N}}$ as follows: for all $S\subseteq N$,
$z_{i}(S):=\begin{cases}\dfrac{1}{pk-k+1}&\mathrm{if\ }i=1\ \mathrm{and\ }\vskip 3.0ptS=N_{1}^{j}\cup N^{c}\ \mathrm{for\ some\ }j\in[pk-k+1],\\ \vskip 3.0pt\dfrac{1}{pk-k+1}&\mathrm{if\ }2\leq i\leq k\ \mathrm{and\ }S=N_{i}^{j}\ \mathrm{for\ some\ }j\in[pk-k+1],\\ 0&\mathrm{otherwise}.\end{cases}$
@@
$(z_{1},\ldots,z_{k})$ is a feasible solution to Mono-LP-Rel.
@@
To show the feasibility of $(z_{1},\ldots,z_{k})$, we show that
$\sum_{\begin{subarray}{c}S\subseteq N,e\in S\\ i\in[k]\end{subarray}}z_{i}(S)=1$ (3.2)
@@
$\sum_{\begin{subarray}{c}S\subseteq N,e\in S\\ i\in[k]\end{subarray}}z_{i}(S)=\sum_{j\in[pk-k+1]}z_{1}(N_{1}^{j}\cup N^{c})=\sum_{j\in[pk-k+1]}\dfrac{1}{pk-k+1}=1.$
@@
$\sum_{\begin{subarray}{c}S\subseteq N,e\in S\\ i\in[k]\end{subarray}}z_{i}(S)$ $=\sum_{\begin{subarray}{c}j\in[pk-k+1]\\ v_{1}\geq j\end{subarray}}z_{1}(N_{1}^{j}\cup N^{c})+\sum_{\begin{subarray}{c}j\in[pk-k+1]\\ 2\leq i\leq k,v_{i}\geq j\end{subarray}}z_{i}(N_{i}^{j})$
@@
The second last equality follows since $\sum_{i=1}^{k}v_{i}=pk-k+1$. Hence, (3.2) holds for every $e\in N$, which implies that $(z_{1},\ldots,z_{k})$ is a feasible solution to Mono-LP-Rel. ∎
By Lemma 3.2 and the optimality of $y^{*}_{1},\ldots,y^{*}_{k}$, we have
$\sum_{S\subseteq N,i\in[k]}y^{*}_{i}(S)f_{i}(S)\leq\sum_{S\subseteq N,i\in[k]}z_{i}(S)f_{i}(S)$
@@
Let $X_{1},\ldots,X_{k}$ be any partition of $N$. Then, we have
$\sum_{i=1}^{k}f_{i}(X_{i})\geq pk+k.$
@@
If $X_{j}\cap Z_{j}\neq\emptyset$ for some $j\in[k]$, then we have
$\sum_{i=1}^{k}f_{i}(X_{i})\geq f_{j}(X_{j})=2pk+1\geq pk+k.$
Hence, it suffices to consider the case when $X_{i}\cap Z_{i}=\emptyset$ for every $i\in[k]$. For each $i\in[k]$, define $x^{i}\in\mathcal{V}$ as follows:
@@
Then, since $X_{i}\cap Z_{i}=\emptyset$ for each $i\in[k]$, we have $e_{x^{i}}\in X_{i}$ for every $i\in[k]$. Hence, we have $X_{i}\cap N_{i}\neq\emptyset$ for every $i\in[k]$. For each $i\in[k]$, we define $b_{i}:=\min\{v_{i}\mid e_{v}\in X_{i}\cap N_{i}\}$. We now show that $\sum_{i=1}^{k}b_{i}\leq pk$. Suppose to the contrary that $\sum_{i=1}^{k}b_{i}\geq pk+1$. Then, since $\sum_{i=1}^{k}(b_{i}-1)\geq pk-k+1$, there exists $v\in\mathcal{V}$ such that $v_{i}\leq b_{i}-1$ for every $i\in[k]$. Choose $i\in[k]$ such that $e_{v}\in X_{i}$. Since $X_{i}\cap Z_{i}=\emptyset$ by the assumption, we have $e_{v}\in X_{i}\cap N_{i}$, which contradicts the definition of $b_{i}$. Hence, we have $\sum_{i=1}^{k}b_{i}\leq pk$. Therefore, since $X_{i}\cap N_{i}\neq\emptyset$ and $X_{i}\cap Z_{i}=\emptyset$ for each $i\in[k]$, we have
$\sum_{i=1}^{k}f_{i}(X_{i})=\sum_{i=1}^{k}\max\{0,\max\{2p+1-v_{i}\mid e_{v}\in X_{i}\cap N_{i}\}\}$
$\geq\sum_{i=1}^{k}\max\{2p+1-v_{i}\mid e_{v}\in X_{i}\cap N_{i}\}=2pk+k-\sum_{i=1}^{k}\min\{v_{i}\mid e_{v}\in X_{i}\cap N_{i}\}$
@@
$\min\left\{\sum_{i=1}^{k}f_{i}(X_{i})\ \middle|\ X_{1},\ldots,X_{k}\ \mathrm{is\ a\ partition\ of\ }N\right\}$
@@
$\leq\min\left\{\sum_{i=1}^{k}f_{i}(X_{i})\ \middle|\ X_{1},\ldots,X_{k}\ \mathrm{is\ a\ partition\ of\ }N\right\}.$

### v2

@@ 2 Approximation algorithm
Since $M$ is divisible by k(k1)k(k-1), we can denote $M=k(k-1)\cdot m$ for some integer $m$. For $k$ positive integers $a_{1},a_{2},\ldots,a_{k}\in[M]$, we say that an ordered tuple $(a_{1},a_{2},\ldots,a_{k})$ covers $N$ if $\bigcup_{i=1}^{k}U_{i}^{a_{i}}=N$. Note that if an ordered tuple $(a_{1},a_{2},\ldots,a_{k})$ covers $N$, then there exist $S_{i}\subseteq U_{i}^{a_{i}}$ for $i=1,2,\ldots,k$ such that $(S_{1},S_{2},\ldots,S_{k})$ is a partition of $N$ and $\sum_{i=1}^{k}f_{i}(S_{i})\leq\sum_{i=1}^{k}f_{i}(U_{i}^{a_{i}})$. We first show that if there exist 2m(k1)2m(k-1) pairwise disjoint ordered tuples covering $N$, then one of them induces a partition $(S_{1},S_{2},\ldots,S_{k})$ of $N$ that corresponds to a $k/2$-approximate solution to $\mathrm{(LP_{mono})}$. We then construct 2m(k1)2m(k-1) such pairwise disjoint ordered tuples covering $N$, which completes the proof of Theorem 1.1. The following lemma is the first part of the proof:
@@
    every ordered tuple covers $N$.
Then, the integrality gap of $\mathrm{(LP_{mono})}$ is at most $k/2$.
@@
$\sum_{S\subseteq N,i\in[k]}z_{i}(S)f_{i}(S)=\sum_{i\in[k],j\in[M]}f_{i}(U_{i}^{j})\geq\sum_{i\in[k],j\in[2m(k-1)]}f_{i}(U_{i}^{a_{i}^{j}})$
$\geq 2m(k-1)\cdot\min\left\{\sum_{i\in[k]}f_{i}(U_{i}^{a_{i}^{j}})\ \middle|\ j\in[2m(k-1)]\right\}.$ (2.2)
@@
$\sum_{S\subseteq N,i\in[k]}z_{i}(S)f_{i}(S)\geq 2m(k-1)\cdot\sum_{i\in[k]}f_{i}(U_{i}^{a_{i}^{j^{*}}}).$
Since $z_{i}=M\cdot y^{*}_{i}=mk(k-1)\cdot y^{*}_{i}$ for each $i\in[k]$, this implies
@@
Since $\bigcup_{i=1}^{k}U_{i}^{a_{i}^{j^{*}}}=N$, there exists a partition $S_{1},S_{2},\ldots,S_{k}$ of $N$ such that $S_{i}\subseteq U_{i}^{a_{i}^{j^{*}}}$ for every $i\in[k]$. Then, by (2.3) and the monotonicity of $f_{i}$, we have
$\dfrac{k}{2}\cdot\sum_{S\subseteq N,i\in[k]}y^{*}_{i}(S)f_{i}(S)\geq\sum_{i\in[k]}f_{i}(U_{i}^{a_{i}^{j^{*}}})\geq\sum_{i\in[k]}f_{i}(S_{i}).$
This implies that the integrality gap of $\mathrm{(LP_{mono})}$ is at most $k/2$. ∎
We now prove Theorem 1.1 by constructing 2m(k1)2m(k-1) pairwise disjoint ordered tuples covering $N$, using only the smallest $2/k$ fraction of the indices $j=1,2,\ldots,M$ of $U_{i}^{j}$ for each $i\in[k]$.
Proof of Theorem 1.1.
By Lemma 2.4, it suffices to show that there exist 2m(k1)2m(k-1) ordered tuples satisfying the conditions in Lemma 2.4. For each positive integer $a\in\mathbb{Z}_{>0}$, define $\mathrm{mod}(a)$ to be the unique integer $r\in[2m(k-1)]$ such that $a\equiv r\pmod{2m(k-1)}$. For each $i\in[k]$ and $j\in[2m(k-1)]$, define $a_{i}^{j}$ as follows:
@@
We show that $(a_{i}^{j})_{i\in[k],j\in[2m(k-1)]}$ satisfies conditions (1)–(3) in Lemma 2.4. For condition (1), we show that $a_{i}^{j}\in[M]$ for each $i\in[k]$ and $j\in[2m(k-1)]$. If $1\leq i\leq k-1$, then we have
@@
The proofs of Lemma 2.4 and Theorem 1.1 provide the construction of a partition $S_{1},S_{2},\ldots,S_{k}$ of $N$ that corresponds to a $k/2$-approximate solution to $\mathrm{(LP_{mono})}$, and hence it is also a $k/2$-approximate solution to MMSCA. The construction works as follows. We first compute a rational optimal solution $(y^{*}_{1},y^{*}_{2},\ldots,y^{*}_{k})$ to $\mathrm{(LP_{mono})}$ whose supports are chains, and define $U_{i}^{j}$ for each $i\in[k]$ and $j\in[M]$. We then define 2m(k1)2m(k-1) ordered tuples $(a_{1}^{j},a_{2}^{j},\ldots,a_{k}^{j})$ covering $N$ as in the proof of Theorem 1.1, and compute the index $j^{*}$ minimizing $\sum_{i=1}^{k}f_{i}(U_{i}^{a_{i}^{j}})$ in the proof of Lemma 2.4. Finally, we obtain sets $U_{1}^{a_{1}^{j^{*}}},U_{2}^{a_{2}^{j^{*}}},\ldots,U_{k}^{a_{k}^{j^{*}}}$, which induces a desired partition $S_{1},S_{2},\ldots,S_{k}$ of $N$.
We prove Theorem 1.2 by showing that the above construction can be done in polynomial time.
Proof of Theorem 1.2.
We can obtain a rational optimal solution $(y_{1}^{*},y_{2}^{*},\ldots,y_{k}^{*})$ to $\mathrm{(LP_{mono})}$ whose supports $\mathcal{L}_{1},\mathcal{L}_{2},\ldots,\mathcal{L}_{k}$ are chains in polynomial time via the ellipsoid method [5] (see Appendix A for details). Then, we define $U_{i}^{j}$ for each $i\in[k]$ and $j\in[M]$, and $a_{i}^{j}$ for each $i\in[k]$ and $j\in[2m(k-1)]$ as in the proof of Theorem 1.1. If we can compute
$j^{*}\in\underset{j\in[2m(k-1)]}{\operatorname{argmin}}\sum_{i=1}^{k}f_{i}(U_{i}^{a_{i}^{j}}),$
then any partition $S_{1},S_{2},\ldots,S_{k}$ of $N$ such that $S_{i}\subseteq U_{i}^{a_{i}^{j^{*}}}$ for every $i\in[k]$ satisfies
$\sum_{i=1}^{k}f_{i}(S_{i})\leq\sum_{i=1}^{k}f_{i}(U_{i}^{a_{i}^{j^{*}}})\leq\dfrac{k}{2}\cdot\sum_{S\subseteq N,i\in[k]}y^{*}_{i}(S)f_{i}(S)$
$\leq\dfrac{k}{2}\cdot\min\left\{\sum_{i=1}^{k}f_{i}(T_{i})\ \middle|\ T_{1},T_{2},\ldots,T_{k}\ \mathrm{is\ a\ partition\ of\ }N\right\},$
as desired. Hence, it suffices to show that $j^{*}$ can be computed in polynomial time. To find $j^{*}$, we set $j_{1}:=1$. Then, for q=2,3,,q=2,3,..., we define $j_{q}$ as the minimum integer such that $j_{q}>j_{q-1}$ and there exists $i\in[k]$ with $U_{i}^{a_{i}^{j_{q}}}\neq U_{i}^{a_{i}^{j_{q-1}}}$. Recall that $n=|N|$. Since for each $i\in[k]$, we have $U_{i}^{1}\supseteq U_{i}^{2}\supseteq\cdots\supseteq U_{i}^{M}$ and there are at most $|\mathcal{L}_{i}|\leq n$ distinct sets among $\{U_{i}^{j}\}_{j=1}^{M}$, we only need to define $j_{q}$ for $q=1,2,\ldots,Q$ with $Q\leq kn$. Since we have
$j^{*}\in\underset{j\in\{j_{1},j_{2},\ldots,j_{Q}\}}{\operatorname{argmin}}\sum_{i\in[k]}f_{i}(U_{i}^{a_{i}^{j}}),$
we can find $j^{*}$ by computing the values of $\sum_{i\in[k]}f_{i}(U_{i}^{a_{i}^{j_{q}}})$ for all $q\in[Q]$. By the definition of $U_{i}^{j}$ and $a_{i}^{j}$, we can compute $j_{q}$ in polynomial time for each $q\in[Q]$. Hence, since $Q\leq kn$, we can compute $j_{q}$ and $\sum_{i\in[k]}f_{i}(U_{i}^{a_{i}^{j_{q}}})$ for all $q\in[Q]$ in polynomial time. ∎
@@ 3 Lower bound on the integrality gap
In this section, we show that the integrality gap of $\mathrm{(LP_{mono})}$ is at least $\frac{k}{2}-\frac{k(k-1)}{\log n}$ when $\log n\geq k$. To this end, we explicitly construct monotone submodular functions $f_{1},f_{2},\ldots,f_{k}$ such that the integrality gap of $\mathrm{(LP_{mono})}$ is at least $\frac{k}{2}-\frac{k(k-1)}{\log n}$.
@@
We now show that for the monotone submodular functions $f_{1},f_{2},\ldots,f_{k}$ defined above, any optimal solution $(y_{1}^{*},\ldots,y_{k}^{*})$ to $\mathrm{(LP_{mono})}$ satisfies
@@
$\leq\min\left\{\sum_{i=1}^{k}f_{i}(S_{i})\ \middle|\ S_{1},S_{2},\ldots,S_{k}\ \mathrm{is\ a\ partition\ of\ }N\right\}.$ (3.1)
The inequality (3) implies that the integrality gap of $\mathrm{(LP_{mono})}$ is at least $\frac{k}{2}-\frac{k(k-1)}{\log n}$, which completes the proof of Theorem 1.3.
To prove (3), we first give an upper bound on the value $\sum_{S\subseteq N,i\in[k]}y^{*}_{i}(S)f_{i}(S)$. To this end, we construct a feasible solution to $\mathrm{(LP_{mono})}$. For each $i\in[k]$ and $j\in[pk-k+1]$, let $N_{i}^{j}:=\{e_{v}\in N\mid v\in\mathcal{V},v_{i}\geq j\}$. Note that $N_{i}^{1}=N_{i}$. In what follows, we will use $N_{i}^{1}$ instead of $N_{i}$ for convenience. Let $N^{c}:=N\setminus\{e_{v}\in N\mid v\in\mathcal{V}\}$. For each $i\in[k]$, define $y^{\prime}_{i}\in\mathbb{R}_{+}^{2^{N}}$ as follows: for all $S\subseteq N$,
$y^{\prime}_{i}(S):=\begin{cases}\dfrac{1}{pk-k+1}&\mathrm{if\ }i=1\ \mathrm{and\ }\vskip 3.0ptS=N_{1}^{j}\cup N^{c}\ \mathrm{for\ some\ }j\in[pk-k+1],\\ \vskip 3.0pt\dfrac{1}{pk-k+1}&\mathrm{if\ }2\leq i\leq k\ \mathrm{and\ }S=N_{i}^{j}\ \mathrm{for\ some\ }j\in[pk-k+1],\\ 0&\mathrm{otherwise}.\end{cases}$
@@
$(y^{\prime}_{1},y^{\prime}_{2},\ldots,y^{\prime}_{k})$ is a feasible solution to $\mathrm{(LP_{mono})}$.
@@
To show the feasibility of $(y^{\prime}_{1},y^{\prime}_{2},\ldots,y^{\prime}_{k})$, we show that
$\sum_{\begin{subarray}{c}S\subseteq N,e\in S\\ i\in[k]\end{subarray}}y^{\prime}_{i}(S)=1$ (3.2)
@@
$\sum_{\begin{subarray}{c}S\subseteq N,e\in S\\ i\in[k]\end{subarray}}y^{\prime}_{i}(S)=\sum_{j\in[pk-k+1]}y^{\prime}_{1}(N_{1}^{j}\cup N^{c})=\sum_{j\in[pk-k+1]}\dfrac{1}{pk-k+1}=1.$
@@
$\sum_{\begin{subarray}{c}S\subseteq N,e\in S\\ i\in[k]\end{subarray}}y^{\prime}_{i}(S)$ $=\sum_{\begin{subarray}{c}j\in[pk-k+1]\\ v_{1}\geq j\end{subarray}}y^{\prime}_{1}(N_{1}^{j}\cup N^{c})+\sum_{\begin{subarray}{c}j\in[pk-k+1]\\ 2\leq i\leq k,v_{i}\geq j\end{subarray}}y^{\prime}_{i}(N_{i}^{j})$
@@
The second last equality follows since $\sum_{i=1}^{k}v_{i}=pk-k+1$. Hence, (3.2) holds for every $e\in N$, which implies that $(y^{\prime}_{1},y^{\prime}_{2},\ldots,y^{\prime}_{k})$ is a feasible solution to $\mathrm{(LP_{mono})}$. ∎
By Lemma 3.2 and the optimality of $y^{*}_{1},y^{*}_{2},\ldots,y^{*}_{k}$, we have
$\sum_{S\subseteq N,i\in[k]}y^{*}_{i}(S)f_{i}(S)\leq\sum_{S\subseteq N,i\in[k]}y^{\prime}_{i}(S)f_{i}(S)$
@@
Let $S_{1},S_{2},\ldots,S_{k}$ be any partition of $N$. Then, we have
$\sum_{i=1}^{k}f_{i}(S_{i})\geq pk+k.$
@@
If $S_{j}\cap Z_{j}\neq\emptyset$ for some $j\in[k]$, then we have
$\sum_{i=1}^{k}f_{i}(S_{i})\geq f_{j}(S_{j})=2pk+1\geq pk+k.$
Hence, it suffices to consider the case when $S_{i}\cap Z_{i}=\emptyset$ for every $i\in[k]$. For each $i\in[k]$, define $x^{i}\in\mathcal{V}$ as follows:
@@
Then, since $S_{i}\cap Z_{i}=\emptyset$ for each $i\in[k]$, we have $e_{x^{i}}\in S_{i}$ for every $i\in[k]$. Hence, we have $S_{i}\cap N_{i}\neq\emptyset$ for every $i\in[k]$. For each $i\in[k]$, we define $b_{i}:=\min\{v_{i}\mid e_{v}\in S_{i}\cap N_{i}\}$. We now show that $\sum_{i=1}^{k}b_{i}\leq pk$. Suppose to the contrary that $\sum_{i=1}^{k}b_{i}\geq pk+1$. Then, since $\sum_{i=1}^{k}(b_{i}-1)\geq pk-k+1$, there exists $v\in\mathcal{V}$ such that $v_{i}\leq b_{i}-1$ for every $i\in[k]$. Choose $i\in[k]$ such that $e_{v}\in S_{i}$. Since $S_{i}\cap Z_{i}=\emptyset$ by the assumption, we have $e_{v}\in S_{i}\cap N_{i}$, which contradicts the definition of $b_{i}$. Hence, we have $\sum_{i=1}^{k}b_{i}\leq pk$. Therefore, since $S_{i}\cap N_{i}\neq\emptyset$ and $S_{i}\cap Z_{i}=\emptyset$ for each $i\in[k]$, we have
$\sum_{i=1}^{k}f_{i}(S_{i})=\sum_{i=1}^{k}\max\{0,\max\{2p+1-v_{i}\mid e_{v}\in S_{i}\cap N_{i}\}\}$
$\geq\sum_{i=1}^{k}\max\{2p+1-v_{i}\mid e_{v}\in S_{i}\cap N_{i}\}=2pk+k-\sum_{i=1}^{k}\min\{v_{i}\mid e_{v}\in S_{i}\cap N_{i}\}$
@@
$\min\left\{\sum_{i=1}^{k}f_{i}(S_{i})\ \middle|\ S_{1},S_{2},\ldots,S_{k}\ \mathrm{is\ a\ partition\ of\ }N\right\}$
@@
$\leq\min\left\{\sum_{i=1}^{k}f_{i}(S_{i})\ \middle|\ S_{1},S_{2},\ldots,S_{k}\ \mathrm{is\ a\ partition\ of\ }N\right\}.$